# Kevin Lin3, Xiaofei Wang3, Zhengyuan Yang3, Linjie Li3, Chung-Ching Lin3, Jianwen Xie4

### v1

@@ 4 Measuring Human Agreement
We conduct human study on edits made by four exemplary models, Step1X-Edit, AnyEdit, Gemini-Flash 2.0 and Flux.1-Kontext-dev, on EdiVal-Bench, generated by EdiVal-Agent as described in Section 3.3. For each edit, we collect two human ratings for each edit, yielding a total of $572\times 4\times 2=4{,}576$ annotations. Raters were recruited online, each holding at least a bachelor’s degree. During evaluation, they were shown the original image, the edited image, and the corresponding instruction. They were asked a binary question: “Evaluate whether the edited image successfully follows the given instruction.”
@@ Results.
Figure 3 summarizes the findings. EdiVal-Agent achieves a human agreement accuracy of 81.3%, significantly higher than VLM-only (Qwen-2.5-VL, 75.2%), CLIP_dir (65.4%), and other zero-shot VLMs. These results verify that integrating VLMs reasoning with object detection leads to better alignment with human judgment compared to existing methods. The inter-annotator’s agreement rate (85.5%) indicates the best performance any evaluation tool can reach.
We attribute the improvement in instruction-following evaluation to two factors. First, for symbolically verifiable instruction types—subject_add, subject_remove, subject_replace, position_change, and count_changeEdiVal-Agent relies solely on Grounding-DINO. It determines the success of an edit by checking object presence/absence, the positions of object centers, and the number of bounding boxes. Results for position_change and subject_remove show that these fixed rules, combined with Grounding-DINO, can significantly outperform Qwen2.5-VL in edit evaluation. We hypothesize that errors in position_change stem from poor spatial reasoning, while failures in subject_remove are due to hallucinations regarding object existence. Second, semantically verifiable types—color_alter, material_alter, text_change, and background_change—are evaluated using Qwen2.5-VL combined with Grounding-DINO. The decomposition stage in EdiVal-Agent can supports evaluation by localizing text regions, enabling the LLM to reason more precisely about text edits. These findings demonstrate that EdiVal-Agent not only enhances interpretability but also improves the practical applicability of evaluation pipelines in real-world scenarios requiring human-like understanding.
@@ 5 Benchmarking Multi-turn Editing
Editing Models
We evaluate three families of image editors: (i) diffusion-based models (InstructPix2Pix (IP2P), MagicBrush, UltraEdit, AnyEdit), (ii) flow-matching models (FLUX.1-Kontext-dev, Qwen-Image-Edit, Step1X-Edit, OmniGen), and (iii) autoregressive/regeneration models (Gemini 2.0 Flash, Nano Banana, and GPT-Image-1) [^1] These models are closed-source and lack public technical reports. We label them “Autoregressive” because they are integrated into an autoregressive language model in the web UI..
EdiVal-Bench
We evaluate all models on EdiVal-Bench, generated by EdiVal-Agent as described in Section 3.3. The source images are from GEdit-Bench Liu et al. (2025), which contains 606 real-world images. For safety, we exclude 34 images with sensitive personal content, yielding 572 images successfully decomposed by EdiVal-Agent. During instruction generation we set MAX_TURNS=3, producing 1,716 editing instructions spanning 9 types (Table 1). In short, every model is evaluated on the same images and the same instructions, using the same metrics.
Metrics
As described in Section 3.4, we evaluate three complementary dimensions. First, instruction following $\alpha_{i}$ is the image-success rate at turn $i$ — the fraction of images for which all edits up to and including turn $i$ are successful. Second, content consistency captures the stability of unchanged objects and background across turns. Consistency is computed primarily using DINOv3 feature similarity for the object and background regions; we also report pixel-level $L_{1}$ distances as diagnostic statistics. Let $\kappa_{i}$ denote the content-consistency score at turn $i$, defined as the mean of the object and background DINOv3 similarities. To summarize overall performance on these two dimensions, we combine with a geometric mean: $\emph{Overall}\;=\;\sqrt{\alpha_{i}\,\kappa_{i}},$ which favors models that achieve balanced gains in both instruction following and content preservation. Perceptual quality (e.g., aesthetics and luminance-based statistics) is reported separately and is not folded into Overall.
We evaluate multi-turn editing across a mix of closed- and open-source models. Unless noted, higher is better; latency is measured in seconds per image (lower is better). We first summarize overall trends and quantitative results (Section 5.1), then analyze instruction following (Section 5.2) and consistency (Section 5.3). We also assess perceptual quality using both learned human-preference scores and low-level statistics (Section 5.4). Finally, because multi-turn instructions can often be compressed into a single prompt, we compare multi-turn versus single-shot “complex” editing (Section 5.5).
5.1 Overall results
@@
Table 2: Results of multi-turn editing. Instruction following, content consistency, and overall across three sequential editing turns. Overall is the geometric mean of instruction following and content consistency. Best per column in dark red; second-best in lighter red. Latency is seconds per image (lower is better).
Latency Instruction Following Content Consistency Overall
Technique Model Date (s/img) T1 T2 T3 T1 T2 T3 T1 T2 T3
Autoregressive Nano Banana 25.08.26 9.70 70.70 50.66 35.35 93.91 90.48 89.48 81.48 67.70 56.24
GPT-Image-1 25.07.16 71.30 73.12 54.89 38.35 81.00 77.78 75.50 76.96 65.34 53.81
Gemini 2.0 Flash 25.02.05 8.20 68.07 45.96 28.42 90.58 85.10 80.88 78.52 62.54 47.94
Flow Matching Qwen-Image-Edit 25.08.04 115.08 72.90 44.06 22.55 84.22 80.52 77.98 78.36 59.56 41.93
Step1X-Edit 25.04.25 20.42 61.89 34.97 17.83 92.76 88.52 85.21 75.77 55.64 38.98
FLUX.1-Kontext-dev 25.06.25 29.21 59.97 32.69 16.61 95.32 92.24 90.22 75.61 54.91 38.71
OmniGen 24.09.11 19.70 54.72 24.48 10.66 93.00 88.42 83.92 71.34 46.52 29.91
Diffusion AnyEdit 24.11.24 3.93 41.07 16.32 7.22 86.42 78.91 70.10 59.58 35.89 22.50
UltraEdit 24.07.07 3.15 51.37 17.70 6.36 86.80 84.50 82.40 66.78 38.67 22.89
MagicBrush 23.06.16 4.08 42.31 15.73 4.90 86.96 81.26 76.86 60.66 35.75 19.41
IP2P 23.12.15 4.09 37.41 10.66 2.80 76.85 68.36 60.30 53.62 26.99 12.99
Summary. Based on the data in Table 2, across three turns, Nano Banana offers the best speed–quality trade-off—highest Overall at T1/T2/T3 (81.48/67.70/56.24) with 9.7 s/img. GPT-Image-1 delivers the strongest instruction following across turns, but its latency (71.3 s/img) [^2] Closed-source latencies are measured in the provider’s hosted web UI; open-source latencies on a single NVIDIA A100 GPU with default settings. and weaker consistency leave it second in Overall; Nano Banana trails GPT-4o by only $\sim$4.2 (T2) and 3.0 (T3) points on instruction following. For consistency, FLUX.1-Kontext-dev leads across turns with Nano Banana close behind, whereas GPT-Image-1 ranks second-to-last—consistent with more regenerative/restyling behavior that can erode pixel/feature stability despite aesthetic gains. Gemini 2.0 Flash is competitive at T1 (second-best Overall) but exhibits a steeper decline by T3. Among open-source systems, Qwen-Image-Edit is strongest at T1 (Overall 78.36) yet degrades rapidly with additional turns, likely due to exposure bias from single-turn training on real images and a short edit-history window that forces the model to operate on its own outputs, more evidence in Section 5.2 and 5.5.
Figure 4: Marginal instruction-following across turns.
@@ 5.2 Instruction Following
We analyze instruction-following ability across nine edit tasks. For a given turn, the marginal task success rate is the proportion of prompts for which the requested edit in that turn is successfully implemented. In contrast, the instruction following score reported in Table 2 corresponds to the multi-turn task success rate at turn ii—the fraction of images for which all edits up to turn $i$ are successful.
Figure 4 shows the evolution of per-turn performance. Models that are AR (autoregressive) and condition on the full editing history are relatively stable across turns: their marginal task success rates change only slightly between turns. By contrast, non-AR models with a very short history window (effectively conditioning only on the previous output) suffer substantial degradation in marginal task success, particularly for flow-matching style models.
A striking example is Qwen-Image-Edit. It is the strongest open-source system at turn 1 (Overall 78.36 vs. 81.48 for Nano Banana) but degrades much faster over subsequent turns. We hypothesize that this is primarily an exposure bias issue: many single-turn edit models are trained to operate on real images and ground-truth inputs, not on their own earlier outputs. When a model must operate on its own previous edits, small distribution mismatches compound across turns and reduce stability—this effect is exacerbated when the model can attend only to a short slice of the history.
This phenomenon suggests that naively applying single-turn edit models to multi-turn editing will amplify exposure bias and lead to rapid failure. To mitigate this, multi-turn training and inference strategies should be considered, for example: training with model-generated rollouts (teacher-student or scheduled sampling) to reduce distribution mismatch, or increasing the history window / using full-history conditioning when feasible.
In summary, autoregressive/full-history architectures exhibit greater stability in multi-turn editing, whereas short-window non-AR models—especially those trained for single-turn flow matching—are more vulnerable to exposure bias. Future work should evaluate the mitigation strategies above and quantify their impact on multi-turn robustness.
Marginal Success Rate Across Instruction Types.
We analyze marginal subtask success rates across turns for different instruction types. The results for GPT-Image-1 are shown in Fig. 5. Other editing models exhibit similar behavior. GPT-Image-1 performs relatively well on attribute-centric tasks such as color_alter and material_alter, but poorly on position_change and count_change, indicating weaknesses in spatial and numerical reasoning, respectively.
Among all subtasks, count_change is the most challenging. Even the best-performing model (GPT-Image-1) achieves a success rate below 25% at turn 1, while most models remain under 5%. We also provide illustrative examples in Figure 6. Notably, all non-AR models fail on this prompt.
Figure 5: Marginal task success rate grouped by task types for GPT-Image-1. It performs relatively well on attribute-centric tasks such as color_alter and material_alter, but poorly on position_change and count_change, indicating weaknesses in spatial and numerical reasoning, respectively. We note that other editing models exhibit similar behavior.
(a) Base Image
(b) GPT-Image-1 (✓)
(c) Nano Banana (✓)
(d) Gemini 2.0 Flash (×)
(e) Qwen-Image-Edit (×)
@@ 5.3 Content Consistency
Beyond successfully implementing editing instructions, multi-turn editing requires attention to consistency. We evaluate two aspects: (i) unchanged-object consistency, which measures whether objects that are not edited up to turn $i$ remain unchanged, and (ii) background consistency, which assesses whether the background remains stable when it is not explicitly modified.
We can evaluate consistency in two ways: 1) extract object features with DINOv3 and compute cosine similarity; 2) measure pixel-level consistency using the $L_{1}$ distance and report one minus this value. Note that when calculating consistency, the grounding box is extracted from the raw input image and applied to all edited images. We therefore choose to report DINOv3 for consistency computation because even small shifts in object location can lead to large variations in pixel-wise $L_{1}$ loss, even when unchanged objects are well preserved. By relying on DINO features, we ensure that consistency is measured semantically, capturing attributes such as object identity, attributes, and texture, etc. Nevertheless, the consistency scores from DINOv3 remain highly correlated with those computed using pixel-wise $L_{1}$ loss (See results in the Appendix).
Based on the results, the most consistent editing model is FLUX.1-Kontext-dev, followed by Nano Banana. In contrast, GPT-Image-1 ranks near the bottom, showing notably poor consistency across turns. Representative qualitative examples are shown in Figure 7 and Figure 8.
@@ 5.4 Visual Quality
(a) Base Image
(b) Qwen-Image-Edit
Figure 8: Illustration of background consistency. Instruction: “Remove brick beige house.” The grounding box is derived from the union of the raw and edited images, excluding the removed object region. This box is used to compute background consistency.
5.4.1 Aesthetics
We quantify aesthetics with HPSv3 Ma et al. (2025a), which we found to generalize reliably to our generated images, whereas alternatives (e.g., RAHF Liang et al. (2024)) underperform in this setting. We do not fold these quality metrics into the aggregate “Overall” score, as preferences differ on whether an edited image should strictly preserve the input style or pursue beautification.
To disentangle these preferences, we report the absolute change in aesthetic score relative to the base image: $\Delta_{i}\;=\;\bigl{|}\mathrm{HPS}_{\text{turn }i}-\mathrm{HPS}_{\text{base}}\bigr{|}.$ Smaller $\Delta$ indicates stronger style fidelity to the base image; larger $\Delta$ reflects greater beautification or stylistic drift. As summarized in Table 3, GPT-Image-1 achieves the highest aesthetic scores across turns and remains stable. Qwen-Image-Edit is the next strongest on absolute HPS. For preserving the base image’s look (small $\Delta$), Gemini 2.0 Flash shows the least drift, with Nano Banana also performing well.
Table 3: Human Preference Scores (HPS), absolute change $\Delta$, and techniques across three refinement turns. dark red denotes the best value in the column (i.e., the least difference in human preference score); lighter red denotes the second-best. For the HPS columns higher values are better (stronger perceived aesthetics). For the $\Delta$ columns smaller values are better (stronger fidelity to the base image).
Technique Model Human Preference Score $\Delta$
@@ 5.4.1 Aesthetics
Autoregressive Nano Banana 4.94 5.12 5.26 0.56 0.73 0.88
@@
Flow Matching Qwen-Image-Edit 5.86 5.72 5.15 1.47 1.34 0.77

### v2

@@ 2.5 Measuring Human Agreement
We conduct human study on edits made by four exemplary models, Step1X-Edit, AnyEdit, Gemini 2.0 Flash and Flux.1-Kontext-dev, on EdiVal-Bench, generated by EdiVal-Agent as described in Sec. 2.3. In total, we collect 4,5764{,}576 human annotations of edits. During evaluation, raters were shown the original image, the edited image, and the corresponding instruction, and asked a binary question: “Evaluate whether the edited image successfully follows the given instruction.”
@@ Results.
Figure 4 summarizes the findings. EdiVal-IF achieves a human agreement accuracy of 81.3%, significantly higher than VLM-only (Qwen-2.5-VL, 75.2%), CLIP_dir (65.4%), and other zero-shot VLMs. These results verify that integrating VLMs reasoning with object detection leads to better alignment with human judgment compared to existing methods. The inter-annotator’s agreement rate (85.5%) indicates the best performance any evaluation tool can reach.
We attribute the improvement in instruction-following evaluation to two factors. First, for symbolically verifiable instruction types—subject_add, subject_remove, subject_replace, position_change, and count_change—EdiVal-IF relies solely on Grounding-DINO. It determines the success of an edit by checking object presence/absence, the positions of object centers, and the number of bounding boxes. Results for position_change and subject_remove show that these fixed rules, combined with Grounding-DINO, can significantly outperform Qwen2.5-VL in edit evaluation. We hypothesize that errors in position_change stem from poor spatial reasoning, while failures in subject_remove are due to hallucinations regarding object existence. Second, semantically verifiable types—color_alter, material_alter, text_change, and background_change—are evaluated using Qwen2.5-VL combined with Grounding-DINO. The decomposition stage in EdiVal-Agent can supports evaluation by localizing text regions, enabling the LLM to reason more precisely about text edits. These findings indicate that EdiVal-IF not only enhances interpretability but also improves the practical applicability of evaluation pipelines in real-world settings that demand human-like understanding. Nonetheless, EdiVal-IF has failure modes, which we document and analyze in Appendix. E.
@@ 3 Benchmarking Multi-turn Editing
Table 3: Results of multi-turn editing. EdiVal-IF, EdiVal-$C$ , and EdiVal-O across three sequential editing turns. Best per column in dark red; second-best in lighter red.
Latency EdiVal-IF EdiVal-$C$ EdiVal-O Rank
Technique Model In-Context Date (s/img) T1 T2 T3 T1 T2 T3 T1 T2 T3
Unknown Seedream 4.0 25.09.10 15.78 75.93 55.58 41.59 92.51 88.03 85.86 83.81 69.95 59.76 1
Nano Banana 25.08.26 9.70 70.70 50.66 35.35 93.91 90.48 89.48 81.48 67.70 56.24 2
GPT-Image-1 25.07.16 71.30 73.12 54.89 38.35 81.00 77.78 75.50 76.96 65.34 53.81 3
Gemini 2.0 Flash 25.02.05 8.20 68.07 45.96 28.42 90.58 85.10 80.88 78.52 62.54 47.94 5
Flow Matching FLUX.1-Kontext-max 25.06.03 10.34 69.49 46.89 31.83 93.93 90.90 88.40 80.79 65.29 53.04 4
Qwen-Image-Edit 25.08.04 115.08 72.90 44.06 22.55 84.22 80.52 77.98 78.36 59.56 41.93 6
Step1X-Edit 25.04.25 20.42 61.89 34.97 17.83 92.76 88.52 85.21 75.77 55.64 38.98 7
FLUX.1-Kontext-dev 25.06.25 29.21 59.97 32.69 16.61 95.32 92.24 90.22 75.61 54.91 38.71 8
OmniGen 24.09.11 19.70 54.72 24.48 10.66 93.00 88.42 83.92 71.34 46.52 29.91 9
Diffusion AnyEdit 24.11.24 3.93 41.07 16.32 7.22 86.42 78.91 70.10 59.58 35.89 22.50 10
UltraEdit 24.07.07 3.15 51.37 17.70 6.36 86.80 84.50 82.40 66.78 38.67 22.89 11
MagicBrush 23.06.16 4.08 42.31 15.73 4.90 86.96 81.26 76.86 60.66 35.75 19.41 12
IP2P 23.12.15 4.09 37.41 10.66 2.80 76.85 68.36 60.30 53.62 26.99 12.99 13
Summary of Results.
Figure 5: Marginal Task Success rate across turns.
Table 3 shows that Seedream 4.0 achieves the strongest overall performance, leading in EdiVal-O across all three turns (83.81/69.95/59.76) with competitive latency (15.8 s/img) [^2] Closed-source latencies are measured in the provider’s hosted web UI; open-source latencies on a single NVIDIA A100 GPU with default settings.. Nano Banana offers the best speed–quality trade-off at 9.7 s/img, ranking second in EdiVal-O and staying close to Seedream 4.0 in both instruction following and consistency. GPT-Image-1 excels in instruction following, but its very high latency (71.3 s/img) and weaker consistency lower its overall score, consistent with more regenerative behavior that prioritizes aesthetics over stability. Among open-source systems, Qwen-Image-Edit performs well initially (EdiVal-O 78.36 at T1) but degrades rapidly with additional turns, likely due to exposure bias as discussed below. We can see that there is a clear gap between the performance of closed-source and open-source systems. With the exception of Qwen-Image-Edit, our model rankings exactly match those reported on the Artificial Analysis leaderboard (rank by human vote) as of September 12, 2025; see Appendix. D.
@@ 3.1 Instruction Following
For a given turn, the marginal task success rate (Eqns. 1 and 2) is the proportion of prompts for which the edit requested at that turn is successfully executed. By contrast, the instruction-following score in Table 3 reports the multi-turn task success rate at turn $i$: the logical AND of the EdiVal-IF scores across all edits along the chain. Figure 5 summarizes per-turn performance. High-ranking models—such as Seedream 4.0, Nano Banana, and FLUX.1-Kontext-max—maintain relatively stable EdiVal-IF across turns, even though Seedream 4.0 and FLUX.1-Kontext-max are not in-context editors (they do not condition on prior prompts or intermediate images). In contrast, several other models exhibit a clear decline in marginal success as the number of turns increases.
A particularly salient case is Qwen-Image-Edit. Although it is the strongest open-source system at turn 1 (EdiVal-O $78.36$ vs. $81.48$ for Nano Banana), its performance degrades more rapidly over subsequent turns. We hypothesize that this stems from exposure bias (Ning et al., 2023; Schmidt, 2019): many single-turn editors are trained to operate on real images and ground-truth inputs rather than on their own previous outputs. When asked to edit their own generations, small distributional mismatches compound across turns, reducing stability; this effect is further aggravated when the model can only attend to a limited history.
Marginal Task Success Rate Across Instruction Types.
We analyze marginal subtask success rates across turns for different instruction types. The results for Nano Banana are shown in Fig. 7. Other editing models exhibit similar behavior. Nano Banana performs relatively well on attribute-centric tasks such as color_alter and material_alter, but poorly on position_change and count_change, indicating weaknesses in spatial and numerical reasoning, respectively.
Figure 6: Marginal task success rate grouped by task types for Nano Banana.
Figure 7: Illustration of background consistency. Instruction: “Remove beige brick house.” The grounding box is the union of all object regions from the raw and edited images.
@@ 3.2 Content Consistency
We evaluate two aspects: (i) unchanged-object consistency (Fig. 8), which measures whether objects that are not edited up to turn $i$ remain unchanged, and (ii) background consistency (Fig. 7), which assesses whether the background remains stable when it is not explicitly modified. When calculating consistency, the grounding box is extracted from the raw input image and applied to all edited images. We therefore choose to report DINOv3 over $L_{1}$ distance for consistency computation because even small shifts in object location can lead to large variations in pixel-wise $L_{1}$ loss, even if unchanged objects are well preserved. By relying on DINO features, we ensure that consistency is measured semantically, capturing attributes such as object identity, attributes, and texture, etc. Nevertheless, the consistency scores from DINOv3 remain highly correlated with those computed using pixel-wise $L_{1}$ loss (See results in the Appendix. J). Based on the results, the most consistent editing model is FLUX.1-Kontext-dev, followed by Nano Banana and FLUX.1-Kontext-max. In contrast, GPT-Image-1 ranks near the bottom, showing notably poor consistency across turns. Representative qualitative examples are shown in Figure 8 and Figure 7.
@@ 3.3 Visual Quality
Besdies EdiVal-VQ, we report the absolute change in VQ score relative to the base image: $\text{EdiVal-VQ}\Delta_{i}\;=\;\bigl|\text{EdiVal-VQ}_{\text{turn }i}-\text{EdiVal-VQ}_{\text{base}}\bigr|.$ Smaller $\Delta$ indicates stronger style fidelity to the base image; larger $\Delta$ reflects greater beautification or stylistic drift. As summarized in Table 4, GPT-Image-1 achieves the highest aesthetic scores across turns and the biggest $\Delta$, indicating a significant stylistic shift (Fig. 3). For preserving the base image’s look (small $\Delta$), Gemini 2.0 Flash shows the least drift, with Nano Banana also performing well. We provide low-level exposure statistics analysis in Appendix. G.
[…]
Technique Model HPS $\Delta$
@@
Unknown Seedream 4.0 5.14 5.15 5.15 0.76 0.77 0.77
Nano Banana 4.94 5.12 5.26 0.56 0.73 0.88
@@
Flow Matching FLUX.1-Kontext-max 5.12 5.07 5.04 0.41 0.49 0.47
Qwen-Image-Edit 5.86 5.72 5.15 1.47 1.34 0.77
@@
Figure 9: Turn-3 instruction following: Multi-turn vs. single-shot complex prompts.
3.4 Multi-turn Editing vs. Complex Editing
We compare two strategies for composing multiple edits. In multi-turn editing, instructions are executed sequentially—apply instruction 1, then apply instruction 2 to the result, and so on. In complex editing, we concatenate $C$ instructions into a single prompt and perform one edit (“complex level” $C$, with $C\in\{1,2,3\}$). Empirically (Fig. 9), when a model does not suffer from exposure bias, multi-turn editing tends to yield higher success rates, consistent with a step-by-step “chain of edits” (analogous to chain-of-thought in reasoning). For instance, Nano Banana benefits from the multi-turn formulation. Conversely, when exposure bias is pronounced, compressing instructions into a single, complex prompt can perform better; see Qwen-Image-Edit in Fig. 9.
4 Conclusion
We introduced EdiVal-Agent, an automated, and interpretable framework for evaluating instruction-based image editing. By leveraging symbolic object decomposition, structured instruction generation, and a hybrid evaluation pipeline integrating both specialist tools and vision-language reasoning models, EdiVal-Agent enables fine-grained, object-centric assessment of modern editing systems. Our design emphasizes transparency, extensibility, and real-world applicability—providing a new standard for evaluating multi-turn, compositional visual editing. We hope EdiVal-Agent will serve as a valuable resource for benchmarking, diagnosing, and advancing the next generation of instruction-based editing models. Limitations and discussions are deferred to Appendix B.
Ethics Statement
Our work focuses on developing reliable and interpretable evaluation methods for instruction-based image editing. While such technology holds promise for creative design, accessibility, and efficient content creation, it may also be misused for harmful purposes such as generating misleading, deceptive, or inappropriate content. We emphasize that our benchmark and evaluation framework are intended solely for advancing research in safe and trustworthy generative AI. To mitigate risks, we build on publicly available datasets, apply safety filters to generated images, and encourage responsible use aligned with ethical standards and community guidelines.
Reproducibility statement
We provide complete prompting templates and pseudo-code in the Appendix, along with implementation details and API links. Comprehensive results, datasets, and evaluation metrics are also documented. To ensure full reproducibility, we will release all code, data, and model checkpoints upon acceptance of this manuscript.
Appendix A Related Work
Instruction-based editing models.
InstructPix2Pix (IP2P) Brooks et al. (2023) introduced a two-stage recipe that converts a text-to-image diffusion model Rombach et al. (2022); Zhang et al. (2025a) into an editor: (i) synthesize paired editing data using Stable Diffusion Rombach et al. (2022) and training-free techniques such as Prompt-to-Prompt Hertz et al. (2023); (ii) fine-tune the diffusion model on these pairs. Subsequent systems—MagicBrush Zhang et al. (2023), UltraEdit Zhao et al. (2024), and AnyEdit Yu et al. (2025)—scale this paradigm to large, fine-grained real-image editing. More recent work (e.g., OmniGen Xiao et al. (2025); Wu et al. (2025b), Step1X-Edit Liu et al. (2025), FLUX.1 Kontext Labs et al. (2025), and Qwen-Image-Edit Wu et al. (2025a), Seedream Gao et al. (2025)) adopts task-aware architectures and increasingly leverages flow matching Liu et al. (2022); Lipman et al. (2022); Zhang et al. (2024).
Editing evaluation.
Recent work relies exclusively on VLMs as interpretable judges—e.g., VIEScore Ku et al. (2023), HQ-Edit Hui et al. (2024), and Complex-Edit Yang et al. (2025)—by querying models such as GPT-4o OpenAI (2025) about specific aspects of an edit. While VLMs offer holistic, language-mediated assessments, they are insufficient on their own: they are notoriously poor at spatial reasoning Zhang et al. (2025b); Cheng et al. (2024); Chen et al. (2024); Qharabagh et al. (2024); Chang et al. (2025) and are prone to object hallucinations in existence, category, attributes, and relations Bai et al. (2024); they have limited sensitivity to pixel-level changes and frequently miss subtle, localized modifications Vo et al. (2025) (e.g., fine structures, small attribute shifts, etc.), which are crucial for evaluating content consistency; they are miscalibrated for artifacts and aesthetics Liang et al. (2024); Xu et al. (2023); Ma et al. (2025), which humans are sensitive to. Our approach, EdiVal-Agent, addresses these gaps by integrating VLM-based reasoning with grounding tools, symbolic, object-centric pixel- and semantic-level tools, and human preference models, yielding a precise and interpretable evaluation of instruction-based editing.
Editing tasks.
Appendix B Limitation and Discussion
Given the object-centric evaluations conducted in this study, several limitations warrant consideration. First, our instruction types are limited to object-centric prompts, which may not capture the full range of creative editing requests typical in real-world scenarios. Future research should explore a broader spectrum of instructions, including those involving stylistic changes or complex narrative elements. Additionally, while our work provides a reliable and comprehensive evaluation framework for multi-turn editing, it does not apply the evaluation results to improve the editing models themselves. A straightforward extension would be to use evaluation scores for Best-of-N selection to improve inference-time performance. Future work could also explore post-training methods such as reinforcement learning, treating the evaluation scores as reward signals.
Appendix C More metrics for human agreement
The most straightforward metric is accuracy. Here, we provide more metrics measuring human agreement: Pearson Linear Correlation Coefficient (PLCC), Cohen’s Kappa Coefficient (Kappa) and F1 scores as shown in Tab. 5. However, we note that for 0/1 predictions, correlation metrics like PLCC and Kappa may be considered not suitable for measuring the agreement with human annotators.
Table 5: Task-specific and overall performance comparison across models. Metrics reported: Pearson Linear Correlation Coefficient (PLCC), Cohen’s Kappa, and F1. Best per column highlighted in bold.
Task Type Model PLCC Kappa F1
subject_add CLIP_dir 0.2110 0.1764 0.6914
Qwen2.5-VL 0.5331 0.5264 0.7893
EdiVal-IF 0.5365 0.5364 0.7786
background_change CLIP_dir -0.0329 -0.0076 0.8745
Qwen2.5-VL 0.5792 0.5686 0.8889
EdiVal-IF 0.5244 0.5157 0.8763
subject_remove CLIP_dir -0.0011 -0.0002 0.6592
Qwen2.5-VL 0.1891 0.1758 0.4896
EdiVal-IF 0.5473 0.5409 0.7837
count_change CLIP_dir 0.0456 0.0142 0.0782
Qwen2.5-VL 0.1998 0.1748 0.2162
EdiVal-IF 0.3431 0.3274 0.3571
material_alter CLIP_dir 0.2086 0.2038 0.4561
Qwen2.5-VL 0.8658 0.8616 0.8971
EdiVal-IF 0.4778 0.4624 0.6364
color_alter CLIP_dir 0.1841 0.0874 0.8542
Qwen2.5-VL 0.8409 0.8407 0.9573
EdiVal-IF 0.7820 0.7744 0.9338
position_change CLIP_dir 0.0996 0.0430 0.3285
Qwen2.5-VL -0.0381 -0.0374 0.1798
EdiVal-IF 0.3907 0.3271 0.5000
text_change CLIP_dir 0.6178 0.6173 0.8063
Qwen2.5-VL 0.7161 0.6947 0.8651
EdiVal-IF 0.7438 0.7347 0.8571
subject_replace CLIP_dir 0.0420 0.0121 0.8219
Qwen2.5-VL 0.6028 0.5994 0.8699
EdiVal-IF 0.5533 0.5429 0.8410
Overall CLIP_dir 0.3186 0.2568 0.6858
Qwen2.5-VL 0.6162 0.6161 0.7922
EdiVal-IF 0.6278 0.6273 0.8030
Appendix D Artificial Analysis Leaderboard
We report the leaderboard from the Artificial Analysis website as of September 12, 2025 (Fig. 10). To ensure a fair comparison, we align on the intersection of models evaluated by both platforms and exclude Qwen-Image-Edit. Among the overlapping systems—Seedream 4.0, Nano Banana (Gemini 2.5 Flash), GPT-Image-1 (GPT-4o), FLUX.1-Kontext-max, and Gemini 2.0 Flash—the relative ordering of human votes on Artificial Analysis matches our EdiVal rankings exactly (Table 6), supporting the accuracy of our methodology.
Table 6: Model rankings on the overlapping set. Relative ranks from Artificial Analysis (human votes) vs. EdiVal (ours) as of Sep 12, 2025.
Model Artificial Analysis (Rank) EdiVal (Rank)
Seedream 4.0 1 1
Nano Banana (Gemini 2.5 Flash) 2 2
GPT-Image-1 (GPT-4o) 3 3
FLUX.1-Kontext-max 4 4
Gemini 2.0 Flash 5 5
Figure 10: Artificial Analysis leaderboard (Sep 12, 2025). Screenshot of the public leaderboard used for comparison in Table 6.
Appendix E Failure Case
We discuss a representative failure mode of our evaluation. The most severe errors arise from false positives in Grounding-DINO, despite its strong open-vocabulary performance. Consider the prompt: “Replace [stone gray railing] with [wooden fence].” As shown in Fig. 11, Seedream 4.0 produces an edit that is visually correct. Our rule for subject_replace declares success if, on the edited image, the source object (stone gray railing) is no longer detected while the target object (wooden fence) is detected. However, Grounding-DINO occasionally reports both the source and target objects in the same region with high confidence, incorrectly suggesting that the source object remains and thereby degrading the measured instruction-following accuracy. Improving the reliability of open-vocabulary detection—particularly reducing false positives—would directly improve the fidelity of our evaluation.
(a) Base image
(b) Seedream 4.0 (edited)
(c) Detected: stone gray railing
(d) Detected: wooden fence
Figure 11: Failure due to detector false positives. Although the edit visually replaces the railing with a wooden fence, Grounding-DINO fires on both “stone gray railing” and “wooden fence” in overlapping regions, causing an incorrect failure in our instruction-following metric.
Appendix F Discussion on single-shot complex editing
Appendix G Discussion on Visual Quality
Technique Model Multi-turn (T3) Complex (C3)
In-Context Nano Banana 35.35 28.14
GPT-Image-1 38.35 28.78
Gemini 2.0 Flash 28.42 21.89
Flow Matching Qwen-Image-Edit 22.55 27.62
Step1X-Edit 17.83 15.73
FLUX.1-Kontext-dev 16.61 19.58
OmniGen 10.66 11.01
Diffusion AnyEdit 7.22 2.80
UltraEdit 6.36 8.22
MagicBrush 4.90 4.55
IP2P 2.80 2.80